1.A nuclear reactor produces 20kg of uranium-235. if the half-life is about 70years, how long will it take to decay to 0.1kg

2.What is the half-life of a 100g sample of nitrogen that decays to 12.5g of nitrogen-16 in 21.6s.

3.A sample of 100g has a half-life of 5years. a. How much is left after 5 years.

4.How much is left after 2 half-lives or after 10years

5.How much is left after 1 half- life

6.The half-life of Molybdenum-99 is 67.0hrs. How much of the 1.000-mg sample of Molybdenum-99 remains after 335hrs?

7.The biological effects of a particular source of radiation depend on several factors. List some of these factors.


1
Expert's answer
2021-03-22T07:37:45-0400

1. ) No = 20Kg N = 0.1Kg , t1/2 = 70 years ,

N/No = (1/2)n

0.1/20 = (1/2)n

1/200 = 1/2n

2n = 200

n = 7.6439

n = t/t1/2

7.6439 = t/70

t = 70×7.6439 = 550.36 years

2. ) No = 100g , N = 12.5g , t = 21.6 s

N/No = (1/2)n

12.5/100 = 1/2n

1/8 = 1/2n

2n = 8 , n = 3

n = t/t1/2

3 = 21.6 /t1/2

t1/2 = 7.2 s

3. ) Half life = 5 years

How much is left after 5 year which is half life = M/2 = 100/2 = 50g

4. ) half life = 5year

How much is left after 2 half life or 10 year = M/2/2 = M/4 = 100 /4 = 25 g

5. ) How much is after 1 half life = M/2 = 100/2 = 50g

6. ) n = t/t1/2 = 335 /67 = 5

N/No = (1/2)n , No = 1 mg , N = ?. , n = 5

N/1 = (1/2)5

N = 1/25 = 1/32 = 0.03125 mg

7. ) biological effects of a particular source of radiation depend on several factors

(a) the high energy of the radiation which causes the damage

(b) the penetrating ability of radiation into human tissue

(c) the ionization ability of the radiation to biomolecules , which disturbs their functions